A 45-foot rope is cut into 2 pieces. Seven less than three times the length of the longer rope piece x is five times the length of the shorter piece, which is (45 − x) ft. Find the length of the longer rope piece x (in ft). (Enter an exact number.)
Answer by macston(5194)

S=shorter piece; L=longer piece
.
S=45-L
L=45-S
.
3L-7=5S
(3L-7)/5=S
.
L=45-S
L=45-((3L-7)/5)
5L=225-(3L-7)
5L=225-3L+7
8L=232
L=29
.
S=45-L=45-29=16
.
ANSWER: The long piece is 29 feet and the short piece is 16 feet.
.
CHECK:
3L-7=5S
3(29)-7=5(16)
87-7=80
80=80
.